### v1

# TF-SASM: Training-free Spatial-aware Sparse Memory for Multi-object Tracking

Abstract
Multi-object tracking (MOT) in computer vision remains a significant challenge, requiring precise localization and continuous tracking of multiple objects in video sequences. This task is crucial for various applications, including action recognition and behavior analysis. Key challenges include occlusion, reidentification, tracking fast-moving objects, and handling camera motion artifacts. Past research has explored tracking-by-detection methods and end-to-end models, with recent attention on tracking-by-attention approaches leveraging transformer architectures. The emergence of data sets that emphasize robust reidentification, such as DanceTrack, has highlighted the need for effective solutions. While memory-based approaches have shown promise, they often suffer from high computational complexity and memory usage. We propose a novel sparse memory approach that selectively stores critical features based on object motion and overlapping awareness, aiming to enhance efficiency while minimizing redundancy. Building upon the MOTRv2 model, a hybrid of tracking-by-attention and tracking-by-detection, we introduce a training-free memory designed to bolster reidentification capabilities and preserve the model’s flexibility. Our memory approach achieves significant improvements over MOTRv2 in the DanceTrack test set, demonstrating a gain of 1.1% in HOTA metrics and 2.1% in IDF1 score.
Index Terms:
Multi-object tracking, Sparse memory approach, Reidentification, Tracking-by-attention
I Introduction
In computer vision, tackling multi-object tracking (MOT) poses a significant challenge. This task involves precisely locating and tracking the temporal continuity of multiple objects within a video sequence [milan2016mot16, dendorfer2020motchallenge]. The resultant trajectories of these objects are invaluable for various applications, spanning from action recognition to behavior analysis. The critical challenges of the MOT problem include occlusion, target rotation, target deformation, blurring, fast-motion objects, and the complexities introduced by camera motion [deepsurvey].
Historically, research efforts in Multiple Object Tracking (MOT) have primarily pursued two distinct directions: tracking-by-detection [sort, deepsort] methodologies and end-to-end models [centertrack, tracktor, memot, memotr]. Tracking-by-detection methods typically rely on object detectors to identify individual objects within each frame, followed by the application of filtering techniques like the Kalman filter to establish correspondence between bounding boxes across the video sequence [sort, ocsort, bytetrack]. Some innovations in this realm have focused on enhancing reidentification by leveraging deep features [deepsort]. Additionally, heuristic approaches, such as camera motion compensation [deepocsort, yi2024ucmctrack], have been proposed to address specific challenges in the tracking domain.
In contrast, end-to-end models represent a more recent paradigm shift, consolidating both detection and association tasks within a unified deep-learning framework. Initially, earlier work often used Convolutional Neural Network (CNN) models for detection, followed by the propagation of detected objects into subsequent frames [tracktor, centertrack]. However, with the introduction of the DETR (DEtection TRansformer) model [detr], which leverages the Transformer architecture [transformer] and stands as a state-of-the-art detector at that time, attention-based tracking approaches [trackformer, transtrack, motr], a subtype of end-to-end models, have emerged, demonstrating promising results in various tracking scenarios.
In recent years, the emergence of datasets like SportsMOT [sportsmot] and DanceTrack [dancetrack] has highlighted the critical need for robust reidentification methods, especially in scenarios featuring visually similar objects. While memory-based approaches [memot, memotr] have shown promise in addressing these challenges, they often suffer from high computational complexity and memory usage, primarily due to storing all features for every frame. This redundancy can be further exacerbated by minor object movements between frames in high-fps video.
Our proposed method offers a streamlined solution: a sparse memory that selectively stores critical features based on object motion and overlapping awareness. This approach leverages the intuition that an object’s degree of deformation correlates directly with its movement, thereby minimizing redundancy and enhancing long-temporal Ouinformation.
To validate our hypothesis, we incorporate our memory mechanism into the MOTRv2 [motrv2] model due to its inherent flexibility. MOTRv2 represents a novel hybrid approach, combining tracking-by-attention and tracking-by-detection methodologies, thus harnessing the strengths of robust detectors and a transformer-like architecture for tracking tasks. This architecture exhibits adaptability across diverse scenarios due to its detector and tracker components separation. By extending this flexibility, we introduce a novel training-free memory mechanism for object tracking, thereby enhancing the association performance of the model.
In summary, our contributions are as follows:
  • Introduce a Spatial-aware Sparse Memory, which selectively captures essential features based on object motion.
  • Propose an Overlapping-aware Feature Selector, which effectively reduces noise from overlapping.
  • Demonstrate the effectiveness of our proposed memory through improvements in key association metrics (2.0% AssA score and 2.1% IDF1 score) comparing with MOTRv2 on the DanceTrack test set, validating its impact on the model’s ability to re-identify objects.
II Related work
II-A Tracking-by-detection
Previous methods in multi-object tracking (MOT) have typically followed a similar pipeline: initially, a detector is trained to localize objects in each frame independently, followed by implementing a matching strategy to associate these detections across the video sequence.
One prevalent matching strategy is exemplified by the Simple Online and Realtime Tracking (SORT) [sort] algorithm, which employs a Kalman filter to predict the next state of an object and then utilizes bipartite matching, facilitated by the Hungarian algorithm, to associate these predictions with new detections. Building upon SORT, DeepSORT [deepsort] incorporates additional object feature matching, thereby improving performance in reidentification tasks.
Another noteworthy approach is BYTETrack [bytetrack], which acknowledges that low-confidence detections may arise due to occlusions or blur, potentially leading to missed tracking if disregarded entirely. BYTETrack leverages these low-confidence detections in a second-round association, effectively reducing false-negative cases.
OC-SORT [ocsort] introduces an innovative update mechanism for the Kalman filter: when an object is rediscovered after a period of missed tracking, the filter’s parameters are adjusted based on the observation. Deep OC-SORT [deepocsort] extends this concept by incorporating adaptive updates of deep features for object reidentification, guided by the detection confidence scores. Furthermore, Deep OC-SORT introduces a camera motion compensation technique with adjustments tailored to observation-centric updates, enhancing robustness in tracking scenarios influenced by camera movements.
II-B End-to-end Approaches
End-to-end object tracking models primarily utilize two approaches: tracking-by-regression and tracking-by-attention. In tracking-by-regression, these methods aim to propagate object detections from one frame to the next, rather than relying on matching detections between consecutive frames as in the traditional tracking-by-detection paradigm. In this approach, the object detector not only provides frame-wise detections but also replaces the data association. For instance, previous endeavors such as Tracktor [tracktor] employ a regression head to determine whether an object exists in the subsequent frame. Alternatively, in CenterTrack [centertrack], the model predicts detections based on object centers encoded as a heatmap in the preceding frame.
When DETR [detr] was proposed for the object detection problem, it introduced an alternative approach: a single model capable of both detection and tracking using Transformer architecture [transformer]. Following this architectural trend, a recent breakthrough is TrackFormer [trackformer], which pioneers object tracking within an end-to-end framework using an attention mechanism in Transformer architecture. Specifically, TrackFormer treats objects from the previous frame as object queries in the subsequent frame, similar to the object detection process employed in the DETR model [detr]. This paradigm shift enables TrackFormer to effectively track objects while harnessing the inherent capabilities of Transformer-based architectures.
Similarly, MOTR [motr] shares an identical concept with TrackFormer. However, while TrackFormer learns within two adjacent frames, MOTR proposes mechanisms for training with multiple frames by aggregating track queries using transformer architecture. In contrast, TransTrack [transtrack] employs an Intersection over Union (IoU) matching mechanism instead of relying solely on the regression head to determine new object appearances or objects from the previous frame.
II-C Hybrid Approaches
Although tracking-by-attention methods have demonstrated efficacy, they encounter a fundamental challenge in adapting to diverse object classes, necessitating training data for both detection and tracking simultaneously. To tackle this challenge, MOTRv2 [motrv2] introduces a novel approach that combines the strengths of tracking-by-detection and tracking-by-attention methods. This is achieved by leveraging a robust pretrained detector tailored to specific object classes and employing an attention mechanism to propagate objects to subsequent frames via cross-attention mechanisms.
However, the original work lacks the necessary memory mechanisms for robust reidentification tasks. Indeed, several studies have showcased the efficacy of memory mechanisms in enhancing reidentification [memotr, memot] within the context of the Multiple Object Tracking (MOT) problem. Unlike Single Object Tracking, which involves a smaller number of stored objects, MOT presents a challenge due to its larger number of objects. Departing from prevalent memory-based approaches that employ deep networks to select and densely store object features at each frame, we emphasize using a sparse memory to mitigate memory usage, particularly for retaining long temporal information. This system selectively stores crucial object features based on deformation and occlusion awareness.
III Method
Figure 1: Our proposed method seamlessly integrates into the MOTRv2 [motrv2] model as a training-free module, preserving MOTRv2’s flexibility for various applications. As depicted in the figure, the pipeline takes proposals generated by YOLOX [yolox2021] and previously tracked objects (detailed in Section III-B) as input for MOTRv2. The Spatial-aware Sparse Memory (SASM), further explained in Figure 2, then processes MOTRv2’s outputs.
In this section, we present our method, which builds on the MOTRv2 [motrv2] model, renowned for its flexibility and superior tracking capabilities. Specifically, we revisit the MOTRv2 architecture and introduce our memory module designed to enhance reidentification performance. The pipeline depicted in Figure 1 illustrates how our memory module takes the embeddings of the object produced by MOTRv2 in each frame and undergoes a filtering process to retain the most important features of the object. These selected features are then propagated to the next frame to facilitate object localization. The filtering process consists of two modules: Spatial-Aware Sparse Memory (SASM) and Overlapping-Aware Feature Selector (OFS), which we will elaborate on in Sections III-C and III-D, respectively.
III-A Problem definition
In the Multiple Object Tracking (MOT) problem, we are presented with a video $\mathcal{V}$ consisting of frames $I_{t}$. The primary objectives involve detecting objects within each frame and tracking a specified set of object classes $\mathcal{C}$. For each class, the algorithm aims to output a collection of bounding boxes corresponding to these objects in each video frame $\mathcal{V}$, along with their respective identities. These detections form tracklets $\mathcal{T}$, sequences of detections associated with the same object throughout the entire video.
III-B Revisiting MOTRv2
End-to-end multiple-object tracking frameworks often suffer from sub-par detection performance compared to tracking-by-detection approaches that rely on standalone object detectors. To address this issue, the authors propose integrating the YOLOX [yolox2021] object detector to provide detection priors for MOTR [motr].
III-B1 Proposal query generation
In this approach, for each frame $t$, YOLOX generates $N_{t}$ proposals $P$, each characterized by center coordinates $(x_{t},y_{t})$, height $h_{t}$, width $w_{t}$, and confidence score $s_{t}$. Initially, a query mechanism similar to the DAB-DETR [liu2022dabdetr] architecture is employed for the first frame, using object coordinates as reference points and a shared learnable embedding as proposal queries. The confidence scores are also integrated into the proposal queries $\mathbf{q}_{t}$ using sine-cosine positional encoding.
III-B2 Proposal propagation
After positional encoding by YOLOX proposals $P_{0}$ in the first frame, proposal queries are further refined by self-attention and interact with image features through deformable attention to produce track queries $q_{0}^{tr}$ and relative offsets $(\Delta x,\Delta y,\Delta w,\Delta h)$. The final prediction $Y_{0}$ is obtained by combining the proposals $P_{0}$ with the predicted offsets.
In subsequent frames, the object queries $\mathbf{q}_{t-1}^{tr}$ and predictions from the previous frame are propagated and concatenated with the proposals from the new frame $P_{t}$. A regression head at the end of the model then determines whether the detected objects are newly appearing or previously tracked objects.
III-B3 Discussion
While object features inherently retain some historical information, prior studies have shown the benefits of incorporating explicit object memory to improve reidentification tasks. To this end, we propose a novel memory integration directly into the MOTRv2 model without requiring fine-tuning. Our experiments reveal that dense memory may struggle to maintain sufficient object information over time in scenarios with a limited number of frames, resulting in redundancy and challenges in matching rapidly evolving object states in subsequent frames. Therefore, we introduce a novel mechanism for selecting object features based on movement and overlapping, enabling the use of smaller memory while preserving longer temporal information.
III-C Spatial-aware Sparse Memory
In this study, we propose a hypothesis linking an object’s deformation to its movement. Accordingly, we utilize the object’s center’s displacement as a metric to determine whether its feature should be stored in memory. The memory pipeline is depicted in Figure 2. The algorithm proceeds through the following steps:
Figure 2: The Spatial-aware Sparse Memory (SASM) takes the output of MOTRv2 as input, which includes both object queries and coordinates. From the object coordinates, the module filters out small-displacement objects to accumulate distance to the next frame, and stores only objects with large displacement, as described in Section III-C. The Overlapping-aware Feature Selector (OFS) plays a crucial role in selecting the best features of each instance since the last memory update, as described in Section III-D.
  1. 1.
    Calculating Moving Distance: In frame $t$, the moving distance $d_{i,t}$ of object $i$ is determined by the Euclidean distance between its centers in frames $t$ and $t-1$. In particular, this distance is computed as the $||\cdot||_{2}$ norm of the vector difference between the centers $\mathbf{c}_{i,t}$ and $\mathbf{c}_{i,t-1}$, respectively, where $\mathbf{c}_{i,t}$ denotes the center coordinates of the bounding box for object $i$ in frame $t$, extracted from $Y_{t}$:
    $d_{i,t}=||\mathbf{c}_{i,t}-\mathbf{c}_{i,t-1}||_{2}$ (1)
  2. 2.
    Thresholding Moving Distance: If the computed moving distance $d_{i,t}$ exceeds a predefined threshold, indicating significant motion, the object’s features are stored in memory. This step ensures that only objects undergoing substantial movement are considered for memory storage.
  3. 3.
    Accumulating Moving Distance: If the moving distance $d_{i,t}$ does not surpass the threshold, it is accumulated with the moving distance of the object in the next frame $t+1$. This accumulation accounts for gradual movement across consecutive frames.
    $d_{i,t+1}=d_{i,t+1}+d_{i,t}$
  4. 4.
    Updating Object Query Feature: The new object query feature $\mathbf{q}_{i,t+1}$ for frame $t+1$ is computed as a weighted average of the previous query feature $\mathbf{q}_{i,t}$ and the average query features $\mathbf{q}_{i,j}$ of the $i$-th object stored in memory. This update mechanism incorporates both the current state of the object and historical information stored in memory, ensuring adaptability to changes in object appearance and position over time.
    $\mathbf{q}_{i,t+1}=\alpha\cdot\mathbf{q}_{i,t}^{tr}+\frac{(1-\alpha)}{m}\sum_{j}\mathbf{q}_{i,j}^{tr}$ (2)
    where $j$ indexes frames in the memory, $m$ is the number of frames in the memory, and $\alpha$ is a hyperparameter determining the bias towards the current feature or the feature in memory.
These steps collectively form a coherent framework for selective features stored in memory, leveraging object movement as a key criterion for memory allocation and ensuring robustness in object tracking across frames. This sparse memory strategy not only addresses the issue of bias towards the latest appearances of objects in feature representation but also facilitates the capture of longer temporal information when compared with dense memory.
III-D Overlapping-aware Feature Selector
Based on our empirical observations, we have noted that object features are often erroneously stored during occlusion or when objects overlap, introducing noise into the stored features. To address this issue, we propose a method for selectively choosing object features to store in memory.
A simplistic approach might involve waiting until the object is in a less obstructed location; however, our experiments reveal that this strategy leads to significant temporal gaps between the last feature stored in memory and the current frame. Therefore, instead of merely storing the object’s latest query feature, we opt to select the object feature from the frame with the lowest Intersection over Union (IoU) with other objects since the last feature update.
This combination not only minimizes the inclusion of erroneous or redundant features but also ensures a smoother continuity in feature representation across frames. By selecting object features based on their spatial relationships with other objects, we aim to enhance our memory storage mechanism’s discriminative power and robustness.
Mathematically, we define the stored feature $\mathbf{\hat{q}}_{i,t}$ at frame $t$ as follows:
$\mathbf{\hat{q}}_{i,t}^{tr}=\begin{cases}\mathbf{q}_{i,t}^{tr}&\text{if }\text{IoU}_{i,t}<\text{IoU}_{i,k}\\ \mathbf{q}_{i,k}^{tr}&\text{otherwise}\end{cases}$ (3)
where $\mathbf{q}_{i,t}^{tr}$ represents the feature of object has index $i$ at frame $t$, and $\mathbf{q}_{i,k}^{tr}$ is the feature of object has index $i$ at the frame $k$ with the highest IoU score. This selection process aims to retain more reliable object features in memory, thereby reducing noise and enhancing tracking accuracy.
As a result, this selective feature storage strategy contributes to the overall efficiency and effectiveness of our tracking system, enabling more accurate and reliable object tracking over extended durations. The pseudo-code of the method is represented in detail in Algorithm 1
0:  $\mathbf{q}_{t}^{tr},\mathbf{q}_{t-1}^{tr}$
1:  Set $O$ with object IDs exists in both $\mathbf{q}_{t}^{tr},\mathbf{q}_{t-1}^{tr}$
2:  Compute $\text{IoU}_{i,t}$
3:  Update $\mathbf{\hat{q}}_{i,t}^{tr}$ as Eq. 3
4:  for $i\in O$ do
5:     Calculate $d_{i,t}$ as Eq. 1
6:     if $d_{i,t}>threshold$ then
7:        Update memory by $\mathbf{\hat{q}}_{i,t}^{tr}$
8:        Update $\mathbf{q_{i,t+1}}$ as Eq. 2
9:     else
10:        Accumulate $d_{i,t+1}$ by $d_{i,t}$
11:        $\mathbf{q}_{i,t+1}=\mathbf{q}_{i,t}^{tr}$
12:     end if
13:  end for
14:  return  $\mathbf{q}_{t+1}$
Algorithm 1 Spatial-aware Sparse Memory
IV Experimental Results
IV-A Dataset and Metric
In this paper, we evaluate our methods on the DanceTrack [dancetrack] dataset. This dataset requires tracking multiple objects with uniform appearance and diverse motion. It comprises 100 videos, with 40 for training (annotations public), 25 for validation (annotations public), and 35 for testing (annotations unpublic). To ensure fair comparisons, we assess our methods using well-known metrics as employed in previous works [motr, motrv2]: HOTA [hota] to evaluate our method and analyze the contribution decomposed into detection accuracy (DetA) and association accuracy (AssA); MOTA [clearmot]; IDF1 [id-metrics] metrics.
IV-B Implementation Details
This study leverages the power of MOTRv2 [motrv2], a model jointly trained on the CrowdHuman [crowdhuman] and DanceTrack [dancetrack] datasets, for object tracking. We utilize pre-trained weights from MOTRv2 and employ YOLOX [yolox2021] for object proposal generation, using the publicly available pre-trained model from the DanceTrack GitHub repository [^1] https://github.com/DanceTrack/DanceTrack. Other MOTRv2 settings remain unchanged.
For the Spatial-aware Sparse Memory (SASM), we utilize two hyperparameters: the maximum number of stored objects (capped at 10) and the moving distance threshold (set to 0.1 of the image size), which were determined through experiments detailed in Section IV-D3.
IV-C Comparing with SOTA
To demonstrate the effectiveness of our proposed method, we compared our performance with existing state-of-the-art (SOTA) methods on the DanceTrack dataset represented in Table I. Our method’s ability to leverage sparse memory for enhanced object re-identification and tracking consistency during object deformation drives this achievement. As shown in Table I, our approach achieves significant improvements over MOTRv2, with gains of 2.0 and 2.1 points, respectively, in AssA and IDF1 metrics. These improvements translate to a HOTA score of 71.2, which convincingly underscores the efficacy of our memory-based strategy. Furthermore, despite employing the same object proposals, our method exhibits a marginal but notable DetA improvement of 0.3 compared to MOTRv2. This enhancement can be attributed to our method’s superior object-tracking consistency.
TABLE I: Performance comparison of SOTA methods.
Methods HOTA DetA AssA MOTA IDF1
FairMOT [fairmot] 39.7 66.7 23.8 82.2 40.8
CenterTrack [centertrack] 41.8 78.1 22.6 86.8 35.7
TransTrack [transtrack] 45.5 75.9 27.5 88.4 45.2
TraDes [trades] 43.3 74.5 25.4 86.2 41.2
ByteTrack [bytetrack] 47.7 71.0 32.1 89.6 53.9
GTR [gtr] 48.0 72.5 31.9 84.7 50.3
QDTrack [qdtrack] 54.2 80.1 36.8 87.7 50.4
MOTR [motr] 54.2 73.5 40.2 79.7 51.5
OC-SORT [ocsort] 55.1 80.3 38.3 92.0 54.6
MeMOTR [memotr] 68.5 80.5 58.4 89.9 71.2
MOTRv2 [motrv2] 69.9 83.0 59.0 91.9 71.7
MOTRv3 [motrv3] 70.4 83.8 59.3 92.9 72.3
Ours 71.2 83.3 61.0 92.0 73.8
IV-D Ablation study
To comprehensively evaluate the efficacy of our proposed method, we conduct experiments across three critical dimensions: assessing the effectiveness of each module, exploring alternative design choices for these modules, and rigorously selecting hyperparameters. To ensure fairness in comparison, we perform an ablation study on the validation set to identify the optimal design choices and hyperparameters.
IV-D1 Effectiveness of each module
To initiate our experimentation, we systematically incorporated each module into the baseline model, MOTRv2 [motrv2], in an iterative manner, leading to the results depicted in Table II. Our analysis indicates that integrating the Spatial-aware Sparse Memory (SASM), as elaborated in Section III-C, improves the IDF1 score by 1.0 and the AssA score by 1.0. Nevertheless, we observed performance degradation in specific videos due to noise stemming from object feature overlap within the stored memory. We introduce the Overlapping-aware Feature Selector (OFS) detailed in Section III-D to tackle this issue. This addition substantially improves tracking consistency by enhancing both detection accuracy (DetA by 0.3) and tracking accuracy (AssA by 0.4).
TABLE II: The effectiveness of each proposed module
Methods HOTA ($\uparrow$) DetA ($\uparrow$) AssA ($\uparrow$) IDF1 ($\uparrow$)
Baseline 65.4 79.0 54.4 67.1
+ SASM 66.1 (+0.6) 79.1 (+0.1) 55.4 (+1.0) 69.4 (+1.0)
+ OFS 66.5 (+0.4) 79.4 (+0.3) 55.8 (+0.4) 69.5 (+0.1)
IV-D2 Alternative design comparison
Our work introduces two memory modules: the Spatial-aware Sparse Memory (SASM) and the Overlapping-aware Feature Selector (OFS). We conducted a series of empirical studies to optimize each module’s design. The first investigation focused on memory density, comparing feature storage at every frame (dense memory) with a sparse approach (Table III). In both experiments, the maximum stored features were capped at 10. In particular, dense memory exhibited significantly lower performance across all metrics in the DanceTrack validation set. We attribute this disparity to rapid object deformation. In particular, for stationary objects, dense memory features become biased towards the current state, leading to missed detections when objects undergo sudden deformations, a frequent occurrence in the DanceTrack dataset.
TABLE III: Comparison between Dense and Sparse memory
Methods HOTA ($\uparrow$) DetA ($\uparrow$) AssA ($\uparrow$) IDF1 ($\uparrow$)
Dense 64.5 78.3 53.3 67.2
Sparse 66.1 (+1.6) 79.1 (+0.8) 55.3 (+2.0) 69.4 (+2.2)
Our next investigation addressed the challenge of handling overlapping object features during storage. As described in Section III-D, two approaches were considered: delaying storage until a frame with a good location of the object is encountered or storing the feature at the best location since the last storage. Our findings revealed that the delay approach introduced a significant drawback: large gaps between feature storage events due to frequent object overlaps (Table IV). Both experiments utilized the same SASM module to ensure a fair comparison.
TABLE IV: Comparison between Delaying and OFS strategy
Methods HOTA ($\uparrow$) DetA ($\uparrow$) AssA ($\uparrow$) IDF1 ($\uparrow$)
Delaying 66.0 79.0 55.3 68.8
OFS 66.5 (+0.5) 79.4 (+0.4) 55.8 (+0.5) 69.5 (+0.7)
Figure 3: Qualitative results for tracking consistency
Figure 4: Qualitative results for tracking re-id performance
IV-D3 Hyperparameters
We conducted thorough experiments to explore our algorithm’s sensitivity to hyperparameters. Initially, we focused on determining the optimal number of frames to store in memory for each object. Our investigation revealed that utilizing 10 frames yielded the most favorable outcomes. Conversely, as detailed in Table V, alternative configurations resulted in notably inferior performance. This discrepancy arises from the inherent challenge of aligning memory features with object features, a task complicated by the absence of explicit training in our algorithm. The more extended number of frames ($\geq 15$ frames) of memory exacerbates this issue by introducing sparse data and noise from tracking failures, leading to a substantial mismatch between memory and object features. Conversely, a paucity of frames (5 frames) constrains the algorithm’s capacity to capture sufficient temporal information.
TABLE V: Comparison between difference memory length
# frames HOTA ($\uparrow$) DetA ($\uparrow$) AssA ($\uparrow$) IDF1 ($\uparrow$)
5 66.0 79.1 55.2 68.9
10 66.5 79.4 55.8 69.5
15 65.8 79.1 54.9 68.5
20 65.6 78.6 54.9 68.6
Furthermore, we conducted experiments to evaluate the impact of adjusting the threshold on the algorithm’s performance as in Table VI. Our analysis indicates that setting the threshold ($\epsilon$) to 0.05 closely aligns the algorithm’s performance with that of Dense memory, as illustrated in Table III. However, surpassing a threshold of 0.1 relative to the image size leads to excessively prolonged movement distances. This delay in storing image features in memory exacerbates the disparity between stored and current object features, resulting in a rapid performance decline hovering around 65.8%. Nevertheless, even with highly sparse memory, the algorithm exhibits performance enhancements compared to the baseline, achieving 65.4%.
TABLE VI: Comparison between difference moving threshold
$\epsilon$ HOTA ($\uparrow$) DetA ($\uparrow$) AssA ($\uparrow$) IDF1 ($\uparrow$)
0.05 65.5 79.1 54.5 68.7
0.1 66.5 79.4 55.8 69.5
0.2 65.7 79.1 54.8 68.5
0.3 65.9 79.2 55.1 68.7
0.4 65.8 79.0 55.0 68.5
The ablation study on hyperparameters underscores a critical issue inherent in the method: its sensitivity to hyperparameter settings. This sensitivity arises from the challenge of effectively managing the gap between memory features and current features.
IV-E Qualitative result
Our approach stands out for its proficiency in tracking objects across two crucial qualitative scenarios: maintaining consistency during changes in viewpoint and re-identifying them after missed detections. As depicted in Figure 3, consider a person (ID 3) whose appearance alters due to a body rotation. In such instances, MOTRv2 faces challenges, often assigning a new ID when the person turns around. Conversely, our method capitalizes on its extensive temporal memory, ensuring the retention of the object’s original ID.
Furthermore, in Figure 4, observe the individual with ID 4, who occasionally experiences mis-detection in varying light conditions. Despite this challenge, our methods maintain the object’s ID even after the model redetects the object in the final frame. In contrast, MOTRv2 not only struggles to associate the misdetected object but also erroneously links the person with ID 3 to the one with ID 4, creating a new ID for the individual previously identified as ID 4 in the preceding frame.
In conclusion, our qualitative analysis highlights the robustness and efficacy of our method in tracking objects in challenging scenarios. From maintaining consistency during viewpoint changes to preserving object IDs after missed detections, our approach demonstrates superior performance to existing models like MOTRv2.
V Conclusion
In conclusion, by introducing a novel memory mechanism, our study builds upon the previous MOTRv2 framework, a hybrid architecture for Multiple Object Tracking (MOT) problems. Specifically, we focus on optimizing feature selection for object memory within the MOT context, which often involves numerous objects and quickly escalates in complexity with dense memory. Unlike prior approaches, we propose a sparse memory scheme based on object movement.
This study presents an effective method for sparsely storing object features based on their movement and overlapping patterns. As demonstrated through experiments, this results in significant performance enhancements over baseline methods. Additionally, we thoroughly evaluate each model design to glean deeper insights into this approach.
However, despite the benefits of our training-free approach, we acknowledge that the sensitivity of memory object features to hyperparameters can pose challenges in aligning them with the MOTRv2 pipeline. This issue warrants further investigation to design adaptors that can seamlessly integrate the features of each module into an end-to-end model architecture.
[…]

### v2

@@
 Multi-object tracking (MOT) in computer vision remains a significant challenge, requiring precise localization and continuous tracking of multiple objects in video sequences.
-This task is crucial for various applications, including action recognition and behavior analysis.
-Key challenges include occlusion, reidentification, tracking fast-moving objects, and handling camera motion artifacts.
-Past research has explored tracking-by-detection methods and end-to-end models, with recent attention on tracking-by-attention approaches leveraging transformer architectures.
 The emergence of data sets that emphasize robust reidentification, such as DanceTrack, has highlighted the need for effective solutions.
-While memory-based approaches have shown promise, they often suffer from high computational complexity and memory usage.
-We propose a novel sparse memory approach that selectively stores critical features based on object motion and overlapping awareness, aiming to enhance efficiency while minimizing redundancy.
-Building upon the MOTRv2 model, a hybrid of tracking-by-attention and tracking-by-detection, we introduce a training-free memory designed to bolster reidentification capabilities and preserve the model’s flexibility.
-Our memory approach achieves significant improvements over MOTRv2 in the DanceTrack test set, demonstrating a gain of 1.1% in HOTA metrics and 2.1% in IDF1 score.
+While memory-based approaches have shown promise, they often suffer from high computational complexity and memory usage due to storing feature at every single frame.
+In this paper, we propose a novel memory-based approach that selectively stores critical features based on object motion and overlapping awareness, aiming to enhance efficiency while minimizing redundancy.
+As a result, our method not only store longer temporal information with limited number of stored features in the memory, but also diversify states of a particular object to enhance the association performance.
+Our approach significantly improves over MOTRv2 in the DanceTrack test set, demonstrating a gain of 2.0% AssA score and 2.1% in IDF1 score.
 Index Terms:
@@
 Tracking-by-detection methods typically rely on object detectors to identify individual objects within each frame, followed by the application of filtering techniques like the Kalman filter to establish correspondence between bounding boxes across the video sequence [sort, ocsort, bytetrack].
-Some innovations in this realm have focused on enhancing reidentification by leveraging deep features [deepsort].
-Additionally, heuristic approaches, such as camera motion compensation [deepocsort, yi2024ucmctrack], have been proposed to address specific challenges in the tracking domain.
+Some innovations in this realm have focused on enhancing reidentification by leveraging deep features [deepsort, deepocsort, qdtrack].
 In contrast, end-to-end models represent a more recent paradigm shift, consolidating both detection and association tasks within a unified deep-learning framework.
-Initially, earlier work often used Convolutional Neural Network (CNN) models for detection, followed by the propagation of detected objects into subsequent frames [tracktor, centertrack].
+Initially, earlier work often used CNN-based architecture, followed by the propagation of detected objects into subsequent frames [tracktor, centertrack].
 However, with the introduction of the DETR (DEtection TRansformer) model [detr], which leverages the Transformer architecture [transformer] and stands as a state-of-the-art detector at that time, attention-based tracking approaches [trackformer, transtrack, motr], a subtype of end-to-end models, have emerged, demonstrating promising results in various tracking scenarios.
@@
 Our proposed method offers a streamlined solution: a sparse memory that selectively stores critical features based on object motion and overlapping awareness.
-This approach leverages the intuition that an object’s degree of deformation correlates directly with its movement, thereby minimizing redundancy and enhancing long-temporal Ouinformation.
+This approach leverages the intuition that an object’s degree of deformation correlates directly with its movement, thereby minimizing redundancy and enhancing long-temporal information.
 To validate our hypothesis, we incorporate our memory mechanism into the MOTRv2 [motrv2] model due to its inherent flexibility.
 MOTRv2 represents a novel hybrid approach, combining tracking-by-attention and tracking-by-detection methodologies, thus harnessing the strengths of robust detectors and a transformer-like architecture for tracking tasks.
-This architecture exhibits adaptability across diverse scenarios due to its detector and tracker components separation.
+This architecture is adaptable across diverse scenarios due to its separation of detector and tracker components.
 By extending this flexibility, we introduce a novel training-free memory mechanism for object tracking, thereby enhancing the association performance of the model.
@@
 II-A Tracking-by-detection
-Previous methods in multi-object tracking (MOT) have typically followed a similar pipeline: initially, a detector is trained to localize objects in each frame independently, followed by implementing a matching strategy to associate these detections across the video sequence.
-One prevalent matching strategy is exemplified by the Simple Online and Realtime Tracking (SORT) [sort] algorithm, which employs a Kalman filter to predict the next state of an object and then utilizes bipartite matching, facilitated by the Hungarian algorithm, to associate these predictions with new detections.
-Building upon SORT, DeepSORT [deepsort] incorporates additional object feature matching, thereby improving performance in reidentification tasks.
-Another noteworthy approach is BYTETrack [bytetrack], which acknowledges that low-confidence detections may arise due to occlusions or blur, potentially leading to missed tracking if disregarded entirely.
-BYTETrack leverages these low-confidence detections in a second-round association, effectively reducing false-negative cases.
-OC-SORT [ocsort] introduces an innovative update mechanism for the Kalman filter: when an object is rediscovered after a period of missed tracking, the filter’s parameters are adjusted based on the observation.
-Deep OC-SORT [deepocsort] extends this concept by incorporating adaptive updates of deep features for object reidentification, guided by the detection confidence scores.
-Furthermore, Deep OC-SORT introduces a camera motion compensation technique with adjustments tailored to observation-centric updates, enhancing robustness in tracking scenarios influenced by camera movements.
+Previous methods in multi-object tracking (MOT) typically involved training a detector to locate objects in each frame individually, followed by linking these detections across the video sequence.
+One effective tracking method is shown in the SORT [sort] algorithm.
+SORT predicts where an object will go next using a Kalman filter and then uses the Hungarian algorithm to link these predictions with new detections.
+DeepSORT [deepsort] builds on SORT by adding more detailed object feature matching, which helps in identifying objects again after they disappear.
+Another notable approach, BYTETrack [bytetrack], handles situations where detections might have low confidence due to obstacles or blurriness.
+BYTETrack deals with these uncertain detections by associating them in a second pass, which reduces cases of wrongly identifying objects as absent.
+OC-SORT [ocsort] introduces an innovative update mechanism for the Kalman filter: when an object reappears after being lost, the filter’s parameters are adjusted based on the new observation.
+Deep OC-SORT [deepocsort] extends this idea by also updating deep features adaptively for object reidentification, guided by the confidence scores of detections.
 II-B End-to-end Approaches
-End-to-end object tracking models primarily utilize two approaches: tracking-by-regression and tracking-by-attention.
-In tracking-by-regression, these methods aim to propagate object detections from one frame to the next, rather than relying on matching detections between consecutive frames as in the traditional tracking-by-detection paradigm.
-In this approach, the object detector not only provides frame-wise detections but also replaces the data association.
-For instance, previous endeavors such as Tracktor [tracktor] employ a regression head to determine whether an object exists in the subsequent frame.
-Alternatively, in CenterTrack [centertrack], the model predicts detections based on object centers encoded as a heatmap in the preceding frame.
-When DETR [detr] was proposed for the object detection problem, it introduced an alternative approach: a single model capable of both detection and tracking using Transformer architecture [transformer].
-Following this architectural trend, a recent breakthrough is TrackFormer [trackformer], which pioneers object tracking within an end-to-end framework using an attention mechanism in Transformer architecture.
-Specifically, TrackFormer treats objects from the previous frame as object queries in the subsequent frame, similar to the object detection process employed in the DETR model [detr].
-This paradigm shift enables TrackFormer to effectively track objects while harnessing the inherent capabilities of Transformer-based architectures.
-Similarly, MOTR [motr] shares an identical concept with TrackFormer.
-However, while TrackFormer learns within two adjacent frames, MOTR proposes mechanisms for training with multiple frames by aggregating track queries using transformer architecture.
-In contrast, TransTrack [transtrack] employs an Intersection over Union (IoU) matching mechanism instead of relying solely on the regression head to determine new object appearances or objects from the previous frame.
-II-C Hybrid Approaches
-Although tracking-by-attention methods have demonstrated efficacy, they encounter a fundamental challenge in adapting to diverse object classes, necessitating training data for both detection and tracking simultaneously.
-To tackle this challenge, MOTRv2 [motrv2] introduces a novel approach that combines the strengths of tracking-by-detection and tracking-by-attention methods.
-This is achieved by leveraging a robust pretrained detector tailored to specific object classes and employing an attention mechanism to propagate objects to subsequent frames via cross-attention mechanisms.
-However, the original work lacks the necessary memory mechanisms for robust reidentification tasks.
-Indeed, several studies have showcased the efficacy of memory mechanisms in enhancing reidentification [memotr, memot] within the context of the Multiple Object Tracking (MOT) problem.
-Unlike Single Object Tracking, which involves a smaller number of stored objects, MOT presents a challenge due to its larger number of objects.
-Departing from prevalent memory-based approaches that employ deep networks to select and densely store object features at each frame, we emphasize using a sparse memory to mitigate memory usage, particularly for retaining long temporal information.
-This system selectively stores crucial object features based on deformation and occlusion awareness.
+End-to-end object tracking models utilize two main approaches: tracking-by-regression and tracking-by-attention.
+Tracking-by-regression propagates object detections from one frame to the next, bypassing traditional matching methods.
+For example, Tracktor [tracktor] uses a regression head to determine object existence in subsequent frames, while CenterTrack [centertrack] predicts detections based on object centers from previous frames.
+With the introduction of DETR [detr], a new approach emerged, integrating detection and tracking using Transformer architecture.
+TrackFormer [trackformer] builds on this, using an attention mechanism in the Transformer to treat objects from previous frames as queries for subsequent frames, enhancing tracking effectiveness.
+Similarly, MOTR [motr] extends this concept by aggregating track queries across multiple frames.
+TransTrack [transtrack] employs an IoU matching mechanism instead of a regression head for determining new object appearances or tracking previous ones.
+II-C Hybrid Approaches and Memory-based Approaches
+Although tracking-by-attention methods are effective, they struggle with adapting to diverse object classes, requiring simultaneous training for detection and tracking.
+MOTRv2 [motrv2] addresses this by combining tracking-by-detection and tracking-by-attention, using a pretrained detector and an attention mechanism to propagate objects across frames.
+Studies have shown that memory mechanisms enhance reidentification [memotr, memot] in MOT.
+Unlike SOT, which deals with fewer objects, MOT involves more objects, making memory management challenging.
+Therefore, instead of storing object features at each frame, we propose a novel approach for memory, which stores object sparsely.
 III Method
 Figure 1: Our proposed method seamlessly integrates into the MOTRv2 [motrv2] model as a training-free module, preserving MOTRv2’s flexibility for various applications.
-As depicted in the figure, the pipeline takes proposals generated by YOLOX [yolox2021] and previously tracked objects (detailed in Section III-B) as input for MOTRv2.
+As depicted in the figure, the pipeline takes proposals generated by YOLOX [yolox2021] and previously tracked objects (detailed in Section III-A) as input for MOTRv2.
 The Spatial-aware Sparse Memory (SASM), further explained in Figure 2, then processes MOTRv2’s outputs.
-In this section, we present our method, which builds on the MOTRv2 [motrv2] model, renowned for its flexibility and superior tracking capabilities.
-Specifically, we revisit the MOTRv2 architecture and introduce our memory module designed to enhance reidentification performance.
-The pipeline depicted in Figure 1 illustrates how our memory module takes the embeddings of the object produced by MOTRv2 in each frame and undergoes a filtering process to retain the most important features of the object.
-These selected features are then propagated to the next frame to facilitate object localization.
-The filtering process consists of two modules: Spatial-Aware Sparse Memory (SASM) and Overlapping-Aware Feature Selector (OFS), which we will elaborate on in Sections III-C and III-D, respectively.
-III-A Problem definition
-In the Multiple Object Tracking (MOT) problem, we are presented with a video $\mathcal{V}$ consisting of frames $I_{t}$.
-The primary objectives involve detecting objects within each frame and tracking a specified set of object classes $\mathcal{C}$.
-For each class, the algorithm aims to output a collection of bounding boxes corresponding to these objects in each video frame $\mathcal{V}$, along with their respective identities.
-These detections form tracklets $\mathcal{T}$, sequences of detections associated with the same object throughout the entire video.
-III-B Revisiting MOTRv2
-End-to-end multiple-object tracking frameworks often suffer from sub-par detection performance compared to tracking-by-detection approaches that rely on standalone object detectors.
-To address this issue, the authors propose integrating the YOLOX [yolox2021] object detector to provide detection priors for MOTR [motr].
-III-B1 Proposal query generation
+In this section, we present our method, which builds on the MOTRv2 [motrv2] model known for its flexibility and superior tracking capabilities.
+We enhance MOTRv2 by introducing a memory module to improve reidentification performance.
+As shown in Figure 1, our memory module filters the object embeddings produced by MOTRv2 in each frame to retain the most important features.
+These selected features are then propagated to the next frame to aid in object localization.
+The filtering process includes two modules: Spatial-Aware Sparse Memory (SASM) and Overlapping-Aware Feature Selector (OFS), detailed in Sections III-B and III-C respectively.
+III-A Revisiting MOTRv2
+Y Zhang et.
+al.
+propose integrating the YOLOX [yolox2021] object detector to provide detection priors for MOTR [motr].
+By providing detection priors, YOLOX reduces the burden on the tracker, which would otherwise be responsible for handling both detection and tracking tasks.
+III-A1 Proposal query generation
 In this approach, for each frame $t$, YOLOX generates $N_{t}$ proposals $P$, each characterized by center coordinates $(x_{t},y_{t})$, height $h_{t}$, width $w_{t}$, and confidence score $s_{t}$.
-Initially, a query mechanism similar to the DAB-DETR [liu2022dabdetr] architecture is employed for the first frame, using object coordinates as reference points and a shared learnable embedding as proposal queries.
-The confidence scores are also integrated into the proposal queries $\mathbf{q}_{t}$ using sine-cosine positional encoding.
-III-B2 Proposal propagation
+Initially, a query mechanism similar to the DAB-DETR [liu2022dabdetr] architecture is employed for the first frame The confidence scores are also integrated into the proposal queries $\mathbf{q}_{t}$ using sine-cosine positional encoding.
+III-A2 Proposal propagation
 After positional encoding by YOLOX proposals $P_{0}$ in the first frame, proposal queries are further refined by self-attention and interact with image features through deformable attention to produce track queries $q_{0}^{tr}$ and relative offsets $(\Delta x,\Delta y,\Delta w,\Delta h)$.
@@
 In subsequent frames, the object queries $\mathbf{q}_{t-1}^{tr}$ and predictions from the previous frame are propagated and concatenated with the proposals from the new frame $P_{t}$.
-A regression head at the end of the model then determines whether the detected objects are newly appearing or previously tracked objects.
-III-B3 Discussion
-While object features inherently retain some historical information, prior studies have shown the benefits of incorporating explicit object memory to improve reidentification tasks.
-To this end, we propose a novel memory integration directly into the MOTRv2 model without requiring fine-tuning.
-Our experiments reveal that dense memory may struggle to maintain sufficient object information over time in scenarios with a limited number of frames, resulting in redundancy and challenges in matching rapidly evolving object states in subsequent frames.
-Therefore, we introduce a novel mechanism for selecting object features based on movement and overlapping, enabling the use of smaller memory while preserving longer temporal information.
-III-C Spatial-aware Sparse Memory
-In this study, we propose a hypothesis linking an object’s deformation to its movement.
-Accordingly, we utilize the object’s center’s displacement as a metric to determine whether its feature should be stored in memory.
-The memory pipeline is depicted in Figure 2.
-The algorithm proceeds through the following steps:
+III-B Spatial-aware Sparse Memory
+In this study, our objective is to optimize memory utilization in video processing by minimizing redundancy in object information storage.
+We propose a sparse storage strategy tailored to capture significant object deformations between frames, thereby avoiding unnecessary data retention during periods of minimal change.
+Acknowledging the diverse deformation characteristics across objects, a uniform storage approach proves inefficient.
+Consequently, we hypothesize a direct relationship between an object’s movement dynamics and its deformation profile in video sequences, motivating an adaptive memory management scheme.
+Our approach leverages the displacement of each object’s centroid as a decisive criterion for selective memory storage.
+By dynamically adjusting storage intervals based on real-time deformation analysis, our method aims to enhance storage efficiency while minimizing redundancy information.
+Figure 2 delineates the architecture of our proposed memory management system, encompassing the following methodological steps:
 Figure 2: The Spatial-aware Sparse Memory (SASM) takes the output of MOTRv2 as input, which includes both object queries and coordinates.
-From the object coordinates, the module filters out small-displacement objects to accumulate distance to the next frame, and stores only objects with large displacement, as described in Section III-C.
-The Overlapping-aware Feature Selector (OFS) plays a crucial role in selecting the best features of each instance since the last memory update, as described in Section III-D.
+From the object coordinates, the module filters out small-displacement objects to accumulate distance to the next frame, and stores only objects with large displacement, as described in Section III-B.
+The Overlapping-aware Feature Selector (OFS) plays a crucial role in selecting the best features of each instance since the last memory update, as described in Section III-C.
   1. 1.
@@
   2. 2.
-    Thresholding Moving Distance: If the computed moving distance $d_{i,t}$ exceeds a predefined threshold, indicating significant motion, the object’s features are stored in memory.
-This step ensures that only objects undergoing substantial movement are considered for memory storage.
-  3. 3.
-    Accumulating Moving Distance: If the moving distance $d_{i,t}$ does not surpass the threshold, it is accumulated with the moving distance of the object in the next frame $t+1$.
+    Accumulating Moving Distance: If the moving distance $d_{i,t}$ does not surpass $\epsilon$, it is accumulated with the moving distance of the object in the next frame $t+1$.
 This accumulation accounts for gradual movement across consecutive frames.
     $d_{i,t+1}=d_{i,t+1}+d_{i,t}$
-  4. 4.
+    Otherwise, if the computed moving distance $d_{i,t}$ exceeds a predefined threshold $\epsilon$, indicating significant motion, the object’s features are stored in memory.
+This step ensures that only objects undergoing substantial movement are considered for memory storage.
+Additionally, with a higher $\epsilon$, longer temporal information of the object can be stored, but it may also lead to the omission of some important features.
+  3. 3.
     Updating Object Query Feature: The new object query feature $\mathbf{q}_{i,t+1}$ for frame $t+1$ is computed as a weighted average of the previous query feature $\mathbf{q}_{i,t}$ and the average query features $\mathbf{q}_{i,j}$ of the $i$-th object stored in memory.
@@
     where $j$ indexes frames in the memory, $m$ is the number of frames in the memory, and $\alpha$ is a hyperparameter determining the bias towards the current feature or the feature in memory.
-These steps collectively form a coherent framework for selective features stored in memory, leveraging object movement as a key criterion for memory allocation and ensuring robustness in object tracking across frames.
-This sparse memory strategy not only addresses the issue of bias towards the latest appearances of objects in feature representation but also facilitates the capture of longer temporal information when compared with dense memory.
-III-D Overlapping-aware Feature Selector
-Based on our empirical observations, we have noted that object features are often erroneously stored during occlusion or when objects overlap, introducing noise into the stored features.
-To address this issue, we propose a method for selectively choosing object features to store in memory.
-A simplistic approach might involve waiting until the object is in a less obstructed location; however, our experiments reveal that this strategy leads to significant temporal gaps between the last feature stored in memory and the current frame.
-Therefore, instead of merely storing the object’s latest query feature, we opt to select the object feature from the frame with the lowest Intersection over Union (IoU) with other objects since the last feature update.
-This combination not only minimizes the inclusion of erroneous or redundant features but also ensures a smoother continuity in feature representation across frames.
-By selecting object features based on their spatial relationships with other objects, we aim to enhance our memory storage mechanism’s discriminative power and robustness.
-Mathematically, we define the stored feature $\mathbf{\hat{q}}_{i,t}$ at frame $t$ as follows:
+These steps collectively form a coherent framework for selective features stored in memory, leveraging object movement as a key criterion for memory allocation and ensuring the diversity of object state in object tracking across frames.
+As a result, this sparse memory strategy not only addresses the issue of bias towards the latest appearances of objects in feature representation but also facilitates the capture of longer temporal information by reducing redundancy when compared with dense memory.
+III-C Overlapping-aware Feature Selector
+Empirical observations reveal that object features are often incorrectly stored during occlusion or overlap, introducing noise.
+To address this, we propose selectively storing object features in memory.
+A simplistic approach of waiting for unobstructed views leads to significant temporal gaps.
+Instead, we select the feature from the frame with the lowest Intersection over Union (IoU) with other objects since the last update.
+This minimizes erroneous features and ensures smoother continuity.
+We define the stored feature $\mathbf{\hat{q}}_{i,t}$ at frame $t$ as:
 $\mathbf{\hat{q}}_{i,t}^{tr}=\begin{cases}\mathbf{q}_{i,t}^{tr}&\text{if }\text{IoU}_{i,t}<\text{IoU}_{i,k}\\ \mathbf{q}_{i,k}^{tr}&\text{otherwise}\end{cases}$ (3)
-where $\mathbf{q}_{i,t}^{tr}$ represents the feature of object has index $i$ at frame $t$, and $\mathbf{q}_{i,k}^{tr}$ is the feature of object has index $i$ at the frame $k$ with the highest IoU score.
-This selection process aims to retain more reliable object features in memory, thereby reducing noise and enhancing tracking accuracy.
-As a result, this selective feature storage strategy contributes to the overall efficiency and effectiveness of our tracking system, enabling more accurate and reliable object tracking over extended durations.
-The pseudo-code of the method is represented in detail in Algorithm 1
-0:  $\mathbf{q}_{t}^{tr},\mathbf{q}_{t-1}^{tr}$
-1:  Set $O$ with object IDs exists in both $\mathbf{q}_{t}^{tr},\mathbf{q}_{t-1}^{tr}$
-2:  Compute $\text{IoU}_{i,t}$
-3:  Update $\mathbf{\hat{q}}_{i,t}^{tr}$ as Eq. 3
-4:  for $i\in O$ do
-5:     Calculate $d_{i,t}$ as Eq. 1
-6:     if $d_{i,t}>threshold$ then
-7:        Update memory by $\mathbf{\hat{q}}_{i,t}^{tr}$
-8:        Update $\mathbf{q_{i,t+1}}$ as Eq. 2
-9:     else
-10:        Accumulate $d_{i,t+1}$ by $d_{i,t}$
-11:        $\mathbf{q}_{i,t+1}=\mathbf{q}_{i,t}^{tr}$
-12:     end if
-13:  end for
-14:  return  $\mathbf{q}_{t+1}$
-Algorithm 1 Spatial-aware Sparse Memory
+where $\mathbf{q}_{i,t}^{tr}$ is the feature of object $i$ at frame $t$, and $\mathbf{q}_{i,k}^{tr}$ is the feature at frame $k$ with the highest IoU.
+This selection process retains more reliable features, reducing noise and e
+As a result, this selective feature storage strategy contributes to our tracking system’s overall efficiency and effectiveness, enabling more accurate and reliable object tracking over extended durations.
 IV Experimental Results
@@
 In this paper, we evaluate our methods on the DanceTrack [dancetrack] dataset.
-This dataset requires tracking multiple objects with uniform appearance and diverse motion.
-It comprises 100 videos, with 40 for training (annotations public), 25 for validation (annotations public), and 35 for testing (annotations unpublic).
-To ensure fair comparisons, we assess our methods using well-known metrics as employed in previous works [motr, motrv2]: HOTA [hota] to evaluate our method and analyze the contribution decomposed into detection accuracy (DetA) and association accuracy (AssA); MOTA [clearmot]; IDF1 [id-metrics] metrics.
+This dataset requires tracking multiple objects with uniform appearance and diverse motion.To ensure fair comparisons, we assess our methods using well-known metrics as employed in previous works [motr, motrv2]: HOTA [hota] to evaluate our method and analyze the contribution decomposed into detection accuracy (DetA) and association accuracy (AssA); MOTA [clearmot]; IDF1 [id-metrics] metrics.
 IV-B Implementation Details
@@
 Methods HOTA DetA AssA MOTA IDF1
-FairMOT [fairmot] 39.7 66.7 23.8 82.2 40.8
 CenterTrack [centertrack] 41.8 78.1 22.6 86.8 35.7
 TransTrack [transtrack] 45.5 75.9 27.5 88.4 45.2
-TraDes [trades] 43.3 74.5 25.4 86.2 41.2
 ByteTrack [bytetrack] 47.7 71.0 32.1 89.6 53.9
-GTR [gtr] 48.0 72.5 31.9 84.7 50.3
 QDTrack [qdtrack] 54.2 80.1 36.8 87.7 50.4
@@
 To initiate our experimentation, we systematically incorporated each module into the baseline model, MOTRv2 [motrv2], in an iterative manner, leading to the results depicted in Table II.
-Our analysis indicates that integrating the Spatial-aware Sparse Memory (SASM), as elaborated in Section III-C, improves the IDF1 score by 1.0 and the AssA score by 1.0.
+Our analysis indicates that integrating the Spatial-aware Sparse Memory (SASM), as elaborated in Section III-B, improves the IDF1 score by 1.0 and the AssA score by 1.0.
 Nevertheless, we observed performance degradation in specific videos due to noise stemming from object feature overlap within the stored memory.
-We introduce the Overlapping-aware Feature Selector (OFS) detailed in Section III-D to tackle this issue.
+To tackle this issue, we introduce the Overlapping-aware Feature Selector (OFS) detailed in Section III-C.
 This addition substantially improves tracking consistency by enhancing both detection accuracy (DetA by 0.3) and tracking accuracy (AssA by 0.4).
@@
 We attribute this disparity to rapid object deformation.
-In particular, for stationary objects, dense memory features become biased towards the current state, leading to missed detections when objects undergo sudden deformations, a frequent occurrence in the DanceTrack dataset.
-TABLE III: Comparison between Dense and Sparse memory
-Methods HOTA ($\uparrow$) DetA ($\uparrow$) AssA ($\uparrow$) IDF1 ($\uparrow$)
+For stationary objects, dense memory features tend to bias towards the current state, often missing sudden deformations common in the DanceTrack dataset.
+In contrast, our method stores fewer features during minor object deformations while preserving longer temporal information, aiding in more accurate object state associations.
+Our next investigation focused on addressing the challenge of handling overlapping object features during storage.
+As described in Section III-C, we explored two approaches: delaying storage until encountering a frame with optimal object positioning or storing the feature at the best location since the last storage.
+Our findings uncovered a significant drawback with the delay approach: large gaps between feature storage events due to frequent object overlaps (Table III).
+Both experiments used the SASM module to ensure a fair comparison.
+TABLE III: Comparison between Memory Types and Different Strategies
+Design Performance Metrics ($\uparrow$)
+HOTA DetA AssA IDF1
 Dense 64.5 78.3 53.3 67.2
-Sparse 66.1 (+1.6) 79.1 (+0.8) 55.3 (+2.0) 69.4 (+2.2)
-Our next investigation addressed the challenge of handling overlapping object features during storage.
-As described in Section III-D, two approaches were considered: delaying storage until a frame with a good location of the object is encountered or storing the feature at the best location since the last storage.
-Our findings revealed that the delay approach introduced a significant drawback: large gaps between feature storage events due to frequent object overlaps (Table IV).
-Both experiments utilized the same SASM module to ensure a fair comparison.
-TABLE IV: Comparison between Delaying and OFS strategy
-Methods HOTA ($\uparrow$) DetA ($\uparrow$) AssA ($\uparrow$) IDF1 ($\uparrow$)
+Sparse (Ours) 66.1 (+1.6) 79.1 (+0.8) 55.4 (+2.1) 69.4 (+2.2)
 Delaying 66.0 79.0 55.3 68.8
-OFS 66.5 (+0.5) 79.4 (+0.4) 55.8 (+0.5) 69.5 (+0.7)
-Figure 3: Qualitative results for tracking consistency
-Figure 4: Qualitative results for tracking re-id performance
+OFS (Ours) 66.5 (+0.5) 79.4 (+0.4) 55.8 (+0.5) 69.5 (+0.7)
 IV-D3 Hyperparameters
@@
 Our investigation revealed that utilizing 10 frames yielded the most favorable outcomes.
-Conversely, as detailed in Table V, alternative configurations resulted in notably inferior performance.
+Conversely, as detailed in Table IV, alternative configurations resulted in notably inferior performance.
 This discrepancy arises from the inherent challenge of aligning memory features with object features, a task complicated by the absence of explicit training in our algorithm.
@@
 Conversely, a paucity of frames (5 frames) constrains the algorithm’s capacity to capture sufficient temporal information.
-TABLE V: Comparison between difference memory length
-# frames HOTA ($\uparrow$) DetA ($\uparrow$) AssA ($\uparrow$) IDF1 ($\uparrow$)
-5 66.0 79.1 55.2 68.9
-10 66.5 79.4 55.8 69.5
-15 65.8 79.1 54.9 68.5
-20 65.6 78.6 54.9 68.6
-Furthermore, we conducted experiments to evaluate the impact of adjusting the threshold on the algorithm’s performance as in Table VI.
-Our analysis indicates that setting the threshold ($\epsilon$) to 0.05 closely aligns the algorithm’s performance with that of Dense memory, as illustrated in Table III.
-However, surpassing a threshold of 0.1 relative to the image size leads to excessively prolonged movement distances.
-This delay in storing image features in memory exacerbates the disparity between stored and current object features, resulting in a rapid performance decline hovering around 65.8%.
-Nevertheless, even with highly sparse memory, the algorithm exhibits performance enhancements compared to the baseline, achieving 65.4%.
-TABLE VI: Comparison between difference moving threshold
-$\epsilon$ HOTA ($\uparrow$) DetA ($\uparrow$) AssA ($\uparrow$) IDF1 ($\uparrow$)
+We also evaluated the impact of adjusting the threshold on our algorithm’s performance (Table IV).
+Setting the threshold ($\epsilon$) to 0.05 aligns closely aligns the algorithm’s performance with the baseline.
+However, thresholds exceeding 0.1 relative to image size result in prolonged movement distances, delaying feature storage in memory.
+This exacerbates disparities between stored and current object features, leading to a performance decline to around 65.8%.
+Nonetheless, even with highly sparse memory, our algorithm shows performance improvements compared to the baseline, achieving 65.4%.
+TABLE IV: Comparison between different parameters: moving threshold vs memory length
+Parameter Performance Metrics ($\uparrow$)
+HOTA DetA AssA IDF1
+Moving Threshold ($\epsilon$)
 0.05 65.5 79.1 54.5 68.7
@@
 0.4 65.8 79.0 55.0 68.5
+Memory Length (# features)
+5 66.0 79.1 55.2 68.9
+10 66.5 79.4 55.8 69.5
+15 65.8 79.1 54.9 68.5
+20 65.6 78.6 54.9 68.6
 The ablation study on hyperparameters underscores a critical issue inherent in the method: its sensitivity to hyperparameter settings.
@@
 Conversely, our method capitalizes on its extensive temporal memory, ensuring the retention of the object’s original ID.
-Furthermore, in Figure 4, observe the individual with ID 4, who occasionally experiences mis-detection in varying light conditions.
-Despite this challenge, our methods maintain the object’s ID even after the model redetects the object in the final frame.
-In contrast, MOTRv2 not only struggles to associate the misdetected object but also erroneously links the person with ID 3 to the one with ID 4, creating a new ID for the individual previously identified as ID 4 in the preceding frame.
-In conclusion, our qualitative analysis highlights the robustness and efficacy of our method in tracking objects in challenging scenarios.
-From maintaining consistency during viewpoint changes to preserving object IDs after missed detections, our approach demonstrates superior performance to existing models like MOTRv2.
+Figure 3: Qualitative results for tracking consistency
 V Conclusion
@@
 This issue warrants further investigation to design adaptors that can seamlessly integrate the features of each module into an end-to-end model architecture.
+Acknowledgements
+This research is funded by University of Science, VNU-HCM under grant number T2023-93.
+Furthermore, Nguyen Quang Thuc was also funded by the Master, PhD Scholarship Programme of Vingroup Innovation Foundation (VINIF), code VINIF.2022.ThS.JVN.10.
 […]